OpenStudy (anonymous):

What is the slope of the line passing through the points (–1, 3) and (4, –7)? A. B. C. D.

pooja195 (pooja195):

\[\LARGE \frac{y_2-y_1}{x_2-x_1}=slope\]

OpenStudy (alexandervonhumboldt2):

use slope formula: \[\frac{ y_1-y_2 }{ x_1-x_2 }\]

OpenStudy (alexandervonhumboldt2):

\[\frac{ 3-(-7) }{ -1-4 }=?\]

OpenStudy (alexandervonhumboldt2):

solve

OpenStudy (anonymous):

A. 2 B.3/4 C. -4/3 D. -2

OpenStudy (alexandervonhumboldt2):

can you solve the eqaution i gave above?

OpenStudy (anonymous):

i think so

OpenStudy (alexandervonhumboldt2):

what did you get then?

OpenStudy (anonymous):

10/-5

OpenStudy (alexandervonhumboldt2):

yes but this can be simplified to -(10/5)=?

OpenStudy (anonymous):

-2

OpenStudy (alexandervonhumboldt2):

yes correct

OpenStudy (anonymous):

thx

OpenStudy (alexandervonhumboldt2):

:)